# LVLM-Intrepret: An Interpretability Tool for Large Vision-Language Models

### v1

@@ 1 Introduction
Recently, large language models (LLM), such as those in the families of GPT [23] and LLaMAsc [39, 40], have demonstrated astounding understanding and reasoning capabilities, as well as the ability to generate output that adheres to human instructions. Building on this ability, many work, such as GPT-4V [24], Qwen-VL [4], Gemini [37], and LLaVA [14], have introduced visual understanding to LLMs. Through the addition of a vision encoder followed by finetuning on multimodal instruction-following data, these prior work have demonstrated large vision-language models (LVLM) that are able to follow human instructions to complete both textual and visual tasks with great aptitude.
LLMs are rapidly surpassing humans in many tasks such as summarization, translation, general question answering, and even creative writing. However, they are still very prone to hallucination, i.e. the fabrication of untrue information [12, 45]. This phenomenon of hallucination is also seen in LVLMs and may even include additional dimensions stemming from the visual modality [42, 50]. With the introduction of LVLMs and their massively increased number of parameters, interpreting and explaining model outputs to mitigate hallucination is becoming an ever rising challenge. In light of the need to understand the reasoning behind model responses, we present an intrepretability tool for large vision-language models: LVLM-Intrepret. The main contributions of the proposed application are the adaptations of multiple interpretability methods to large vision-language models for interactive analysis. These methods include raw attention, relevancy maps, and causal interpretation. LVLM-Interpret is applicable to any LVLM with a transformer-based LLM front-end. In this paper, we demonstrate a case study of how our application can be used to better understand the inner workings and failure mechanisms in LLaVA [14].
@@
Figure 2: Visualization of crossmodal attentions
@@ 2.1 Interpretability of Vision Models
Studying the intepretability of Vision Transformers (ViT) has gained popularity with task-specific analysis like image captioning [8, 35, 10], object detection [44, 9, 3], image recognition [19, 46]. Recently, there has been an increased demand for interpretability analysis for the medical domain in applications such as pathology [22, 13], retinal image classification [26, 11] and COVID-19 analysis [20, 34] among others. A novel vision transformer was presented in [27] with a training procedure which has an interpretability-aware training objective. [21] proposed a method to use the activations of a vision transformer’s hidden layers to predict the relevant parts of the input that contribute to its final predictions, while [17] introduced quantification indicators to measure the impact of patch interactions to effectively exploit responsive fields of each patch in ViT.
@@ 3 Interface and Interpretability Functions
LVLM-Intrepret was developed using Gradio [1] and follows a standard layout for multimodal chat. Figure 1 shows an example of the user interface. In the UI, a user is able to upload an image and issue multimodal queries to the LVLM. An added editing feature allows for basic modification of the input image to probe the model with adversarial variations. As the LVLM model generates a response, the attention weights of the model are stored internally and are later presented to the user for visualization. The application also constructs relevancy maps and causal graphs relating to the output of the model. Once a response is returned, the user is able to utilize these results as a way to interpret the model output. The following sections describe each of these interpretability functions.
@@ 3.2 Relevancy Map
We adapted the calculation of relevancy maps to LVLMs such as LLaVA. Relevancy scores are backward propagated through the LLM as well as vision transformer commonly used as the vision encoder. For image analysis, the relevancy scores corresponding to image patches are reshaped into a grid that matches the layout of the original image. This grid forms the basis of the relevancy map. The relevancy map is then upscaled to the original image size using bilinear interpolation, providing a visual representation of the parts of the image most relevant to each generated token.
@@ 3.3 Causal Interpretation
Recently, a causal interpretation of the attention mechanism in transformers was presented [33]. This interpretation leads to a method for deriving causal explanations from attention in neural networks (CLEANN). In this sense, if explanation tokens would have been masked in the input, the model would have generated a different output. Such explanations, which are a subset of the input tokens, are generally tangible and meaningful to humans. Here, in addition to the explanation, we also plot the causal graphs around the explained tokens, and allow the user to decrease or increase the explanation set size based on this graph. We refer the reader to [32, 33] for more details.
@@
Consider the following example. A sequence of image token is given as part of the prompt, and the text prompt is ‘What color is the man’s shirt?’. In response, the model generates ‘The man’s shirt is yellow.’. One may be interested in understanding which image tokens are responsible for the the token yellow. That is, identify the parts of the image such that if masked, will cause the model to output a different color. First, the top-$k$ tokens having the highest attention values for yellow are assigned to nodes. Then, using the full attention matrix of the last (deepest) layer is used to learn a causal graph. This causal graph is a partial ancestral graph [30, 47], where a circle edge-mark indicates a non identifiable edge mark (head and tail are equally valid). From this graph, a tree rooted at node yellow is extracted (Figure 4) such that it includes all paths that potentially influence the root [33, Appendix B. Definition 2]. CLEANN searches for the minimal explaining set by gradually increasing the search distance in this tree (radius in Figure 4) from the explained node. We enable setting the search distance manually for examining the output. An example is given in Figure 3. In Figure 3(a), the 50 tokens having the highest attention values are marked. In Figure 3(b)Figure 3(e), tokens within different search distances from the explained token are marked.
This explanation approach solely relies on attention values in the last layer. While raw attention values describe pair-wise, marginal dependence relations between a token and the explained token, a causal-discovery algorithm identifies conditional independence relations. Thus, based only on the current trained weights, it can identify, those tokens that if perturbed may change the generated token.
@@ 4 Case Study
To demonstrate the functionalities of LVLM-Interpret, we analyze the LLaVA model on the Multimodal Visual Patterns (MMVP) benchmark dataset [38]. MMVP focuses on identifying “CLIP-blind pairs”, i.e. images that are perceived as similar by CLIP [28] despite having clear visual differences. This dataset is of particular interest for our study since it highlights the challenges faced in answering relatively straightforward questions, often leading to incorrect responses.
@@ 5 Conclusions and Future Directions
In this paper we presented LVLM-Interpret, an interactive tool for interpreting responses from large vision-language models. The tool offers a way to visualize how generated outputs relate to the input image through raw attention, relevancy maps, and causal interpretation. Through the many interpretability functions, users can explore the inner mechanisms of LVLMs and obtain insights on failure cases. Future work can include consolidation of the multiple methods for a more comprehensive metric to explain the reasoning behind model responses.

### v3

# LVLM-Interpret: An Interpretability Tool for Large Vision-Language Models

@@ 1 Introduction
Recently, large language models (LLM), such as those in the families of GPT [24] and LLaMA [39, 40], have demonstrated astounding understanding and reasoning capabilities, as well as the ability to generate output that adheres to human instructions. Building on this ability, many work, such as GPT-4V [25], Qwen-VL [4], Gemini [37], and LLaVA [14], have introduced visual understanding to LLMs. Through the addition of a vision encoder followed by finetuning on multimodal instruction-following data, these prior work have demonstrated large vision-language models (LVLM) that are able to follow human instructions to complete both textual and visual tasks with great aptitude.
LLMs are rapidly surpassing humans in many tasks such as summarization, translation, general question answering, and even creative writing. However, they are still very prone to hallucination, i.e. the fabrication of untrue information [12, 45]. This phenomenon of hallucination is also seen in LVLMs and may even include additional dimensions stemming from the visual modality [42, 50]. With the introduction of LVLMs and their massively increased number of parameters, interpreting and explaining model outputs to mitigate hallucination is becoming an ever rising challenge. In light of the need to understand the reasoning behind model responses, we present an interpretability tool for large vision-language models: LVLM-Interpret. The proposed application adapts multiple interpretability methods to large vision-language models for interactive analysis. These methods include raw attention, relevancy maps, and causal interpretation. LVLM-Interpret is applicable to any LVLM with a transformer-based LLM front-end. We further demonstrate how we can gain insight on the inner workings of LVLMs using our application.
The main contributions of this paper are:
  • We propose an interactive tool for interpreting the inner attention mechanisms of large vision-language models
  • We present a case study that sheds light on a possible cause behind certain failure cases in LVLMs
  • Through a study on causal explanations, we postulate that large vision-language models (such as LLaVA [14]) implicitly learn to represent causal structure
@@
Figure 2: Visualization of cross-modal attentions
@@ 2.1 Interpretability of Vision Models
Studying the interpretability of Vision Transformers (ViT) has gained popularity with task-specific analysis like image captioning [8, 35, 10], object detection [44, 9, 3], image recognition [19, 46]. Recently, there has been an increased demand for interpretability analysis for the medical domain in applications such as pathology [22, 13], retinal image classification [27, 11] and COVID-19 analysis [20, 34] among others. A novel vision transformer was presented in [28] with a training procedure which has an interpretability-aware training objective. [21] proposed a method to use the activations of ViT’s hidden layers to predict the relevant parts of the input that contribute to its final predictions, while [17] introduced quantification indicators to measure the impact of patch interactions to effectively exploit responsive fields of patches in ViT.
@@ 3 Interface and Interpretability Functions
LVLM-Interpret was developed using Gradio [1] and follows a standard layout for multimodal chat. Figure 1 shows an example of the user interface. In the UI, a user is able to upload an image and issue multimodal queries to the LVLM. An added editing feature allows for basic modification of the input image to probe the model with adversarial variations. As the LVLM model generates a response, the attention weights of the model are stored internally and are later presented to the user for visualization. The application also constructs relevancy maps and causal graphs relating to the output of the model. Once a response is returned, the user is able to utilize these results as a way to interpret the model output. The following sections describe each of these interpretability functions.
@@ 3.2 Relevancy Map
We adapted the calculation of relevancy maps to LVLMs such as LLaVA. Relevancy scores are backward propagated through the LLM as well as vision transformer commonly used as the vision encoder. For image analysis, the relevancy scores corresponding to image patches are reshaped into a grid that matches the layout of the original image. This grid forms the basis of the relevancy map. The relevancy map is then upscaled to the original image size using bilinear interpolation, providing a visualization of the regions of the image most relevant to each generated token.
@@ 3.3 Causal Interpretation
Recently, a causal interpretation of the attention mechanism in transformers was presented [33]. This interpretation leads to a method for deriving causal explanations from attention in neural networks (CLEANN). In this sense, if explanation tokens would have been masked in the input, the model would have generated a different output. Such explanations, which are a subset of the input tokens, are generally tangible and meaningful to humans. The method was previously demonstrated for a single modality, such as recommendation systems [23], and text sentiment classification [33]. Here, we enable examining if multi-modal LLMs, which are significantly larger, internally represent causal structures. Hence, in addition to providing causal explanation, we plot the causal graphs around the explained tokens, and allow the user to decrease or increase the explanation set size based on this graph. We refer the reader to [32, 33] for more details.
@@
Consider the following example. A sequence of image token is given as part of the prompt, and the text prompt is ‘What color is the man’s shirt?’. In response, the model generates ‘The man’s shirt is yellow.’. One may be interested in understanding which image tokens are responsible for the the token yellow. That is, identify the parts of the image such that if masked, will cause the model to output a different color. First, the top-k tokens having the highest attention values for yellow are assigned to nodes. Then, using the full attention matrix of the last (deepest) layer is used to learn a causal graph. This causal graph is a partial ancestral graph [30, 47], where a circle edge-mark indicates a non identifiable edge mark (head and tail are equally valid). From this graph, a tree rooted at node yellow is extracted (Figure 4) such that it includes all paths that potentially influence the root [33, Appendix B. Definition 2]. CLEANN searches for the minimal explaining set by gradually increasing the search distance in this tree (radius in Figure 4) from the explained node. An example is given in Figure 3. In Figure 3(a), the 50 tokens having the highest attention values are marked. In Figure 3(b)Figure 3(e), tokens within different search distances from the explained token are marked.
This explanation approach solely relies on attention values in the last layer. While raw attention values describe pair-wise, marginal dependence relations, the causal-discovery algorithm in CLEANN identifies conditional independence relations. Thus, based only on the current trained weights, it can identify those tokens that if perturbed may change the generated token.
@@ 4 Case Study
To demonstrate the functionalities of LVLM-Interpret, we analyze the LLaVA model on samples from the Multimodal Visual Patterns (MMVP) benchmark dataset [38]. MMVP focuses on identifying “CLIP-blind” images that are demonstrably hard for LVLMs to reason on. This dataset is of particular interest for our study since it highlights the challenges faced in answering relatively straightforward questions, often leading to incorrect responses.
@@ 5 Conclusions and Future Directions
In this paper we presented LVLM-Interpret, an interactive tool for interpreting responses from large vision-language models. The tool offers a way to visualize how generated outputs relate to the input image through raw attention, relevancy maps, and causal interpretation. Through the many interpretability functions, users can explore the inner mechanisms of LVLMs and obtain insights on failure cases. The application can also reveal several potential paths for enhancing the performance of LVLMs. Future work can include consolidation of the multiple interpretability methods for a more comprehensive metric to explain the reasoning behind model responses.